2. A tank ( not rigid) contains 2.5 litres of helium gas at 280C . At constant pressure, what will


be the volume of the tank after heating it and its content to 430C temperature?



1
Expert's answer
2022-05-10T07:47:05-0400

V1 / T1 = V2 / T2

2.5 / 280 = V2 / 430

V2 = 3.8 L